Air at atmospheric pressure (1.0x10^5Pa) is compressed at constant temperature in a bicycle pump so that its volume is reduced by a factor f of 7.

What is the pressure of the air when compressed?
1
Expert's answer
2019-11-28T08:29:00-0500

Use Boyle's Law to find final pressure


"P_1V_1 =P_2V_2"

As

"P_1 = 1.0\times 10^5 Pa"

"V_2 = \frac{V_1}{7}"

find "P_2"



"1.0\times 10^5 V_1 = \frac {V_1}{7} P_2"

"P_2 = 7.0\times 10^5 Pa"